# Quality or Quantity? On Data Scale and Diversity in Adapting Large Language Models for Low-Resource Translation

### v1

@@ 4.7 Discussion: LLMs vs NMT models
For the indigenous American languages, the current SOTA systems are NordicAlps (Attieh et al., 2024) and DC-DMV (DeGenaro and Lupicki, 2024) that report average chrF++ scores of 26.73 and 23.76 respectively. Both are NMT models. NordicAlps trained a multilingual model from scratch on varying mixtures of eng-spa and spa-x data at scale, with their most major gains (+4 chrF++ points) coming from a novel redundancy-driven tokenization method. DeGenaro and Lupicki (2024) fine-tune NLLB-200 using a variety of parallel data sources, but unlike us, they generate a large amount of synthetic data. Our best-performing system, Llama3 with CPT on parallel data and 5 epochs of SFT, yields a score of 20.93 in its most optimal hyperparameter setting. While this is a gain of +17.06 chrF++ from a 5-shot prompted Llama3, it is also almost 6 chrF++ points behind the SOTA. Since ours was not a shared task effort, we did not attempt explorations with tokenization or synthetic data – which are concurrent to our findings and would likely boost our performance further. Overall, our results suggest that low-resourced LLM-MT systems, while promising, are still behind the curve compared to SOTA NMT models.
@@ 5 Conclusion
In this paper, we approach low-resource LLM-MT from a data-centric perspective, and study performance along two axes: i) scale of parallel data (quantity), and ii) diversity in SFT (quality). Through elaborate experiments and analyses with 3 LLMs and 2 language groups, we conclude that quantity plays the dominant role in downstream performance. Specifically, we note that parallel data drives performance singificantly during both CPT and SFT, with HRLs and LRLs displaying different behaviours – making bitext a critical resource even in the modern era of LLM-MT. Moreover, we establish that diversity (on multiple fronts) consistently causes decline in MT quality, with multilingual fine-tuning on task-specific data remaining the most effective option – reaffirming our previous findings on the value of scale. We hope these findings will be useful considerations when scaling to massively multilingual LLMs of the future.
6 Limitations
@@ Appendix A Data
A.1 Data for CPT
We provide the detailed monolingual data statistics for the indigenous American languages used for CPT in Table 10. We collect this data from various sources. For the indigenous American languages, we use MADLAD-400 (Kudugunta et al., 2024), GLOT 500 (ImaniGooghari et al., 2023), Wikipedia, data curated by the University of Helsinki (De Gibert et al., 2023) and OCR data collected by Iyer et al. (2024). The English and Spanish data used for replay comes from Wikipedia and MADLAD-400 too. For the Indic languages, we use MADLAD-400 and GLOT 500 once again, along with the verified split of Sangraha (Khan et al., 2024) – a large corpus for Indian languages.
A.2 Data for SFT
We provide the templates used for generating Machine Translation and XQA Instructions in Tables 11 and 12 respectively. For MT, we randomly use one of these prompts to create an instruction for doing SFT of our model. For XQA, the provided instructions are used to prompt Mixtral-8x-7B to generate a question, which is later used for creating the synthetic XQA datasets - as described in Section 3.2.
@@ A.2 Data for SFT
Aymara 1M 23.4M Quechua 1.9M 45.1M Guarani 1.1M 37.6M
Nahuatl 1.1M 32.1M Otomi 0.6M 23.6M English 0.4M 9.8M
Spanish 0.7M 27.9M Shipibo-Konibo 0.1M 3.3M Bribri 0.1M 2.6M
Asháninka 0.2M 2.2M Chatino 0.3M 5.4M Huichol 0.2M 3.5M
Tarahumara 0.2M 2.3M Total (Replay) 1.1M 37.6M Total 7.8M 218.9M
@@ B.1 Parallel vs Monolingual Data Ratio
In Table 14, we find that mixing higher ratios of parallel data with monolingual data either performs comparably or improves performance. We do not go higher than 50% since our parallel data runs out at this stage, and to ensure higher ratios we would have to oversample the existing dataset – which would not lead to a fair comparison with the other baselines. Nevertheless, given the monotonic trend, it would be interesting to explore if mixing higher ratios of parallel data continues to improve the performance even more. However, it is likely that there is some ceiling as to how much parallel data one should mix – for instance, Alves et al. (2024) show how a baseline trained on 100% parallel data underperforms compared to mixing it with monolingual data.
@@ B.2 Scaling up SFT: Mistral 7B
Figure 7 shows the scaling behaviour of the Mistral 7B model with different pretraining recipes (no CPT, CPT with monolingual data, and CPT with a mixture of monolingual and parallel data) – where terminology is the same as that defined in Section 3.1. We observe that the trends for Mistral are largely similar to Llama3:
@@
    ‘No CPT’ underperforms CPT baseline, but they become comparable at about ~500K sentences. As hypothesised in Section 4.4, due to a 4x larger vocabulary, Llama3 has 4x more parameters to fine-tune [^4] LoRA module parameters are negligible in comparison. For fine-tuning input and output embeddings, Llama3 has 128K * 4096 * 2 $\approx$ 1B parameters, whereas Mistral has 32K * 4096 * 2 $\approx$ 250M parameters. LoRA parameters are on the order of 3M, for comparison. – meaning our Llama3 models are likely more data hungry and retain the benefits of CPT over longer periods, but smaller models overfit sooner – leading to a shorter ‘cross-over’ threshold.
@@
    For HRLs, the performance plateaus quite quickly, while for LRLs SFT quality continues to grow with scale - once again following Llama3’s trend
@@
(a) Average (overall)
(b) Average (high-resource)
(c) Average (low-resourceη)
Figure 7: Scaling up Mistral 7B models with different CPT recipes (no CPT, CPT with monolingual data, and CPT with a mixture of monolingual and parallel data) on MT data for the American languages. We observe that the trends are very similar to Llama3 – with the sole exception that the gains of CPT diminish faster, at around 500K sentences. This might be because Mistral, due to a 4x smaller vocabulary, gets fine-tuned effectively with less data. ‘High-resource’ refers to the relatively higher-resourced languages in our low-resource setup (Aymara, Guarani and Quechua) while the other 8 are grouped as low-resourceη.
@@ B.3 Bilingual vs Multilingual LLM-MT: Detailed Results
We provide detailed, language-specific results for bilingual vs multilingual LLM-MT in Table 13. We observe that multilingual LLM-MT mostly outperforms bilingual baselines – with the exception of 3 relatively higher-resourced languages (Aymara, Nahuatl and Guarani) where there is a bit of a gap.

### v2

@@ 4.7 Discussion: LLMs vs NMT models
For the Indigenous American languages, the current SOTA systems are NordicAlps (Attieh et al., 2024) and DC-DMV (DeGenaro and Lupicki, 2024) that report average chrF++ scores of 26.73 and 23.76 respectively. Both are NMT models. NordicAlps trained a multilingual model from scratch on varying mixtures of eng-spa and spa-x data at scale, with their most major gains (+4 chrF++ points) coming from a novel redundancy-driven tokenization method. DeGenaro and Lupicki (2024) fine-tune NLLB-200 using a variety of parallel data sources, but unlike us, they generate a large amount of synthetic data. Our best-performing model, Llama3 with CPT on parallel data and 5 epochs of SFT, yields a score of 20.93 in the best setting. While this is a gain of +17.06 chrF++ from a 5-shot prompted Llama3, it is also almost 6 chrF++ points behind the SOTA. Since ours was not a shared task effort, we did not attempt explorations with tokenization or synthetic data, which are concurrent to our findings and would likely boost performance further. Overall, our results suggest that low-resourced LLM-MT systems, while promising, are still behind the curve compared to SOTA NMT models.
@@ 5 Conclusion
In this paper, we approach low-resource LLM-MT from a data-centric perspective and study performance along two axes: i) the size of parallel data and ii) diversity in SFT tasks. Through experiments and analyses, we conclude that quantity plays a dominant role in downstream performance. Specifically, parallel data drives performance significantly during both CPT and SFT, with HRLs and LRLs displaying different behaviours—making bitext a critical resource even in the modern era of LLM-MT. Moreover, we establish that diversity (on multiple fronts) consistently declines MT quality, with multilingual fine-tuning on task-specific data being the most effective option, reaffirming our previous findings on the value of scale. We hope these findings will be useful considerations when scaling to massively multilingual LLMs of the future.
Limitations
@@
Acknowledgments
The authors have received funding from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant numbers 10052546 and 10039436].
Computations described in this research were supported by the Baskerville Tier 2 HPC service (https://www.baskerville.ac.uk). Baskerville was funded by the EPSRC and UKRI through the World Class Labs scheme (EP/T022221/1) and the Digital Research Infrastructure programme (EP/W032244/1) and is operated by Advanced Research Computing at the University of Birmingham.
@@ Appendix A Data
A.1 Monolingual Data
We provide the detailed monolingual data statistics for the indigenous American languages used for CPT in Table 10. We collect this data from various sources. For the indigenous American languages, we use MADLAD-400 (Kudugunta et al., 2024), GLOT 500 (ImaniGooghari et al., 2023), Wikipedia, data curated by the University of Helsinki (De Gibert et al., 2023) and OCR data collected by Iyer et al. (2024). The English and Spanish data used for replay comes from Wikipedia and MADLAD-400 too. We use MADLAD-400 and GLOT 500 again for the Indic languages again, along with the verified split of Sangraha (Khan et al., 2024)—a large corpus for Indian languages.
A.2 Parallel Data
We provide the templates used for generating Machine Translation and XQA Instructions in Tables 11 and 12 respectively. For MT, we randomly use one of these prompts to create an instruction for doing SFT of our model. For XQA, the provided instructions are used to prompt Mixtral-8x-7B to generate a question, which is later used for creating the synthetic XQA datasets as described in Section 3.2.
We source the parallel data for generating these from the work of Iyer et al. (2024), which includes data sourced from AmericasNLP’23 official training datasets (Ebrahimi et al., 2024), NLLB-200 and FLORES-200 (Costa-jussà et al., 2022), OPUS (Tiedemann, 2009) and data curated by Helsinki in their 2023 submission (De Gibert et al., 2023).
A.2.1 Cleaning Parallel Data
Rule-based Filtering
We clean parallel data by following standard filtering rules. First, we remove pairs with more digits or non-alphanumeric characters than alphabetic ones on either side. Then, we remove pairs where either the source or the target has less than 3 words or more than 120 words. To adjust for pairs containing partial or incomplete translations, we apply a combination of two filtering rules: a) for sentences with less than 25 words, the character length difference must be less than 65, and b) for those longer than 25 words, the character length ratio between either source and target, or vice versa, must be at most 1.55. Lastly, we filter our sentences with non-Latin characters in either sentence, as well as pairs with identical source and target sentences to create our final training dataset.
Neural Quality Estimation
We also experimented with using neural quality estimation using models like LASER-3 (Heffernan et al., 2022). LASER-3 supports 3 of the 11 American languages—the HRLs (Aymara, Guarani and Quechua)—and we used it for scoring parallel sentences for these languages. We then sorted them in order of decreasing quality, calculated as the cosine similarity of LASER representations. For the other 8 LRLs, we used random scoring to simulate standard shuffling behaviour. We found that training on this ‘sorted’ data actually performed worse than our default baseline trained on unscored corpora, consistently at various SFT scales. As none of the authors spoke these languages, we could not substantively explain why, but we hypothesize that LASER representations might not be very reliable for these very LRLs, and might introduce certain undesirable biases in the sorted data, making the default baseline more robust to different kinds of bitext. In practice, we found that the standard rule-based filtering approaches worked best and thus, we stuck to them for cleaning our data.
Finally, we note that both our monolingual and parallel corpora span a variety of domains similar to our test data. Also, due to the paucity of data, we use all available sources.
@@ A.2.1 Cleaning Parallel Data
Aymara (aym) 1M 23.4M Quechua (quy) 1.9M 45.1M Guarani (grn) 1.1M 37.6M
Nahuatl (nhe) 1.1M 32.1M Otomi (oto) 0.6M 23.6M English (eng) 0.4M 9.8M
Spanish (spa) 0.7M 27.9M Shipibo-Konibo (shp) 0.1M 3.3M Bribri (bzd) 0.1M 2.6M
Asháninka (cni) 0.2M 2.2M Chatino (ctp) 0.3M 5.4M Huichol (hch) 0.2M 3.5M
Tarahumara (tar) 0.2M 2.3M Total (Replay) 1.1M 37.6M Total 7.8M 218.9M
@@ B.1 Parallel vs Monolingual Data Ratio
In Table 14, we find that mixing higher ratios of parallel data with monolingual data either performs comparably or improves performance. We do not go higher than 50% since our parallel data runs out at this stage, and to ensure higher ratios we would have to oversample the existing dataset—which would not lead to a fair comparison with the other baselines. Nevertheless, given the monotonic trend, it would be interesting to explore if mixing higher ratios of parallel data continues to improve the performance even more. However, it is likely that there is some ceiling as to how much parallel data one should mix. For instance, Alves et al. (2024) show how a baseline trained on 100% parallel data underperforms compared to mixing it with monolingual data.
@@ B.2 Scaling up SFT: Mistral 7B
(a) Average (overall)
(b) Average (high-resource)
(c) Average (low-resourceη)
Figure 7: Scaling up Mistral 7B models with different CPT recipes (no CPT, CPT with monolingual data, and CPT with a mixture of monolingual and parallel data) on MT data for the American languages. We observe that the trends are very similar to Llama3, with the sole exception that the gains of CPT diminish faster at around 500K sentences. This might be because Mistral, due to a 4x smaller vocabulary, gets fine-tuned effectively with less data. ‘High-resource’ refers to the relatively higher-resourced languages in our low-resource setup (Aymara, Guarani and Quechua) while the other 8 are grouped as low-resourceη.
Figure 7 shows the scaling behaviour of the Mistral 7B model with different pre-training recipes (no CPT, CPT with monolingual data, and CPT with a mixture of monolingual and parallel data), where terminology is the same as that defined in Section 3.1. We observe that the trends for Mistral are largely similar to Llama3:
@@
    ‘No CPT’ underperforms the CPT baseline, but they become comparable at about ~500K sentences. As hypothesised in Section 4.4, due to a 4x larger vocabulary, Llama3 has 4x more parameters to fine-tune [^4] LoRA module parameters are negligible in comparison. For fine-tuning input and output embeddings, Llama3 has 128K * 4096 * 2 $\approx$ 1B parameters, whereas Mistral has 32K * 4096 * 2 $\approx$ 250M parameters. LoRA parameters are on the order of 3M, for comparison., meaning our Llama3 models are likely more data hungry and retain the benefits of CPT over longer periods, but smaller models overfit sooner, leading to a shorter ‘cross-over’ threshold.
@@
    For HRLs, the performance plateaus quite quickly, while for LRLs SFT quality continues to grow with scale, once again following Llama3’s trend
@@ B.3 Bilingual vs Multilingual LLM-MT: Detailed Results
We provide detailed, language-specific results for bilingual vs multilingual LLM-MT in Table 13. We observe that multilingual LLM-MT mostly outperforms bilingual baselines with the exception of 3 relatively higher-resourced languages (Aymara, Nahuatl and Guarani) where there is a bit of a gap.